# Adaptive Text Watermark for Large Language Models

### v1

@@ 5.5 Ablation Study
In this section, our investigation focuses on assessing the robustness and text quality of watermarked text using our watermark under varying entropy thresholds and watermark strength. Specifically, for this analysis, we adjust the threshold value $\alpha$ to be 0, 1, 2, 3 and set the watermark strength $\delta$ at levels of $0.5$, $1$, $1.5$, and $2$.
Entropy Thresholds. In Table 3, we provide the ROC-AUC, Best F1 Score under different entropy thresholds when paraphrasing is conducted using GPT-3.5. In addition, we provide the Average Watermark Rate (AWR), which is the fraction of the watermarked tokens. Figure 4 shows the text perplexity comparison under different entropy thresholds. When $\alpha=0$, AWTI is skipped in watermark generation. The ROC-AUC and Best F1 Score are high, but the watermarked text exhibits high perplexity, with a median value exceeding $10$. When setting $\alpha$ to 1, 2, or 3, there is a notable decrease in the perplexity of the watermarked text, as fewer tokens are perturbed reflected by the decreasing AWR. In addition, the robustness of the watermark is improved when AWTI is applied ($\alpha=1,2$), potentially because some modified tokens are filtered out by AWTI in detection. For entropy threshold $\alpha=3$, our method can achieve results comparable to other methods, with only about half of the watermarked tokens.
@@
Figure 4: Comparison of text perplexity at varying entropy thresholds and across different watermark strengths. $\alpha$ represents the entropy threshold. $\delta$ represents the watermark strength.
Watermark Strength. In Table 4 and Figure 4, we increase the watermark strength $\delta$ from $0.5$ to $2$ to examine its impact on robustness and perplexity. As shown in Table 4, there is little impact on the ROC-AUC values in terms of robustness. Interestingly, in contrast to previous works (Kirchenbauer et al., 2023a; Zhao et al., 2023) where perplexity tends to rise with larger watermark strength, the perplexity of the text watermarked with our method decreases as the $\delta$ increases in a reasonable range. This benefits from the proposed AWTS, which embeds the watermark by adjusting temperature as explained in Section 3.3. However, this does not imply that a larger $\delta$ consistently results in better text quality. A higher watermark strength can cause the token selection process to resemble greedy sampling more closely, potentially leading to the generation of repetitive content.
@@ 6 Conclusion
In this paper, we propose a holistic LLM watermark method to achieve the following properties: strong robustness, high security, high text quality, and accurate detection without the need to access the original language model and prompt. Our experiments show that, compared to baseline methods, our approach demonstrates better performance under different paraphrase attacks. Notably, our method has a negligible impact on text perplexity when compared to un-watermarked text. Additionally, we highlight the security of our approach by employing the spoofing attack. Specifically, the capacity of language models differs among various models. Therefore, the model owner may need to tune the entropy threshold and watermark strength carefully based on the performance of the language model to get better watermarking performance. Moreover, the watermarked model owners are suggested to train their own small-sized Measurement Model based on the watermarked model for performance and security considerations. In future research, we aim to conduct a theoretical analysis to deepen the understanding of the performance characteristics of the watermarking approach and offer an optimal solution.
Appendix A Additional Experiment Results
A.1 Semantic Mapping Model Training Details
@@
The Semantic Mapping Model is a neural network that starts and ends with linear layers and contains two residual blocks in the middle. For the loss function provided in Section 3.2, the weight of each term is set as 1. For the training process, the optimization algorithm we utilize is stochastic gradient descent.
@@
When we train the Semantic Mapping Model (detailed in Section 3.2), in order to ensure the logits scaling vector is sufficiently resilient to variations of semantics, we rescale the Euclidean distance between two sentence embeddings $d=D(u(S),u(S^{\prime}))$ to a wider range. The original range $d\in[L,U]$, where $L=0$ and $U$ is the maximum distance. To enlarge this range, we apply a linear transformation $(T:\mathbb{R}\rightarrow\mathbb{R})$ to the original distance $d$, resulting in a new distance $d^{\prime}=T(d)\in[L^{\prime},U^{\prime}]$ where $L^{\prime}<L$ and $U^{\prime}>U$. The linear transformation is expressed as
$T(d)=\left(\frac{d-L}{U-L}\right)\cdot(U^{\prime}-L^{\prime})+L^{\prime},$
where we set the $L^{\prime}=-2$ and $U^{\prime}=4$ in our experiment.
A.2 Additional Robustness Results
@@
Figure 5: Comparisons of ROC curves of different watermark methods applied to various language models, assessing their performance against two different paraphrasing attacks.
A.3 Additional Ablation Study
@@
Measurement Models. We investigate how different Measurement Models (MMs) impact the robustness of our watermarks and the quality of the generated text. For this purpose, we employ MMs with different sizes and types, such as GPT2 (124M), GPT2-large (774M), OPT-125M, and OPT-350M, in our evaluation. For all models, we set the $\alpha$ and $\delta$ as $2$ and $3$, respectively. As shown in Table 5, there is only a minor influence on the robustness of the watermark when different MMs are adopted.
However, the effect on text quality is more nuanced. As indicated in Figure 6, models like GPT2-large and OPT-350M will result in lower perplexity compared to the GPT2 and OPT-125M. One primary reason is that different MMs produce different entropy estimates, which results in different AWRs even with the same entropy threshold. Generally, a higher AWR implies a larger number of watermarked tokens, which can marginally impact text quality.
@@
A.4 Security Analysis Details
In this subsection, we provide a detailed introduction to the security analysis outlined in Section 5.4. The security of a watermark primarily relies on the difficulty of decryption by potential attackers. This paper uses the spoofing attack (Sadasivan et al., 2023) to decrypt the watermarking scheme. The spoofing attack aims to uncover the potential ‘green’ token of the watermarking scheme by analyzing the 181 common token frequencies in the watermarked text.
For KGW-0, which uses a fixed ‘green/red’ token list throughout the generation, we directly calculate the frequency of tokens in the watermarked text. In the case of KGW-1/KGW-2, the previous one/two token is used as a hash key to generate a ‘green’ list. Therefore, we examine the occurrences of token pairs in the watermarked text to ascertain the ‘green’ list corresponding to each specific token. In our method, the logits scaling vector varies according to the semantics. We use L3 level data (Musical class) in the DBPedia Classes Dataset as the prompt to generate the watermarked text. For all watermarking methods, we identify the top 50 highest frequency words from the 181 common tokens and calculate the percentage of these words that are actually ‘green’ tokens. For each watermarking method, we produce 5000 instances of watermarked texts, each consisting of 200 tokens.
Appendix B Watermarked Text Generation time
This section compares the watermarked text generation time and the watermark detection time for different approaches. For each method, we generate and detect 100 instances of watermarked text, each containing roughly 200 tokens. Subsequently, we compute the average time taken for both the generation and detection of each instance of watermarked text. For our method, we set the value of the entropy threshold $\alpha=2$ and the watermark strength $\delta=1.5$. We evaluate the generation and detection time of the proposed adaptive watermarking with two different Measurement models.
@@ Appendix B Watermarked Text Generation time
Appendix C Watermarked Text Example
We present some examples of watermarked text using our proposed method. Table 8 shows the un-watermarked text and watermarked text under different entropy thresholds ($\alpha$). The watermarked texts are generated using the OPT-6.7B model with the watermark strength $\delta=1.5$. The watermarked tokens are highlighted in yellow. As the value of $\alpha$ increases, it is evident that the proportion of watermarked tokens noticeably decreases.
@@ Appendix C Watermarked Text Example
It’s not always easy to keep mentally focused during running. Focus requires discipline and a certain level of mental toughness, meaning that you have developed the capacity to perform to the best of your ability despite external conditions or internal distractions.
Un-watermarked text
While it’s not impossible to train for a specific event and keep mentally focused at all times, it’s challenging. Here are three things you can do to help yourself.
Plan Ahead
@@
The best time to think about race strategy is during warm-ups before any race. Once your legs get moving, they will tell you what they need—you just have to listen to them! It’s also a good idea to write down in detail the training plan for your event and make sure that you understand it completely. Plan your warm-up so that there is an emphasis on building speed without injury risk.
Get Out of Your Head
There are lots of other things to focus your attention on besides running. If you’re running around town, take a walk break every quarter mile or so. This allows you to stretch, cool off and rest, while keeping your muscles engaged.
Watermarked text ($\alpha=1$)
@@
The ability and desire for discipline are developed through practice. The first step in developing discipline during running, especially if a runner wants a higher level performance, begins with setting goals and working towards meeting these goals. Goal setting allows runners an avenue to create the mental toughness needed to achieve their best results during training runs or races. Setting realistic expectations will also allow runners to work within their abilities instead of trying to perform beyond what they are currently capable of. Setting realistic expectations can also serve as motivation for runners to continue training despite the obstacles that life has to offer during the course of training or race day. It can also serve as motivation to continue training even when injuries or other problems make training challenging or downright unpleasant. For runners who want to set realistic expectations for themselves during training or race day, the below checklist can serve as guidelines:
Mentally prepared - Have enough sleep the night before the race or training runs
Stress free environment - Remove negative people from the environment
Eat right - Eat healthy meals the day before the race or training runs
Watermarked text ($\alpha=2$)
@@
The best runners are often described in terms like, focused, determined, and driven. You may have read that somewhere on a running magazine cover, but did anyone actually write that down? I was recently asked if I knew any specific exercises to help people stay mentally focused during training runs. My response wasNo,” but the question got me thinking more about how to develop mental toughness for runners. Staying mentally prepared is an important part of the training process that can be just as important as physical preparation, but many runners don’t give it much thought. It’s easy to see how physical preparation is important for race day performance, but what exactly does it mean when you talk about mental preparation? Here are five things you need to do to stay mentally prepared while running:
1. Have realistic goals and expectations. Some people set unrealistic expectations for themselves or others regarding the number of miles they’ll run each week, their times for certain races, or even their overall goal for their training program.
Watermarked text ($\alpha=3$)
@@
The best runners are focused on a specific task, which allows for a level playing field. The same applies in any aspect in which concentration and discipline are necessary. In order for a runner, cyclist, tennis star, golffer, etc. to perform their best, they must have the ability to stay focused at all times. To achieve this state of focus or mental toughness requires training (mental conditioning) and practice. You do not have to run every day for hours, but just like anything else it’s about putting in the time and practicing until you get better over time.
So how does this apply to your life? The best runners know that running is an important part of their lives, as well as their training routine and their overall health. The best runners also understand what is required to keep them mentally strong during the race. They know how to train themselves so that they will always be ready and prepared for the next race, regardless of external conditions or the obstacles within themselves.

### v2

@@ 5.5 Ablation Study
In this section, our investigation focuses on assessing the robustness and text quality of watermarked text using our watermark under varying entropy thresholds and watermark strength. Specifically, we adjust the threshold value $\alpha$ to 0, 1, 2, 3 and set the watermark strength $\delta$ at $0.5$, $1$, $1.5$, and $2$. In Appendix E, we further analyze the performance of our watermark under different measurement models, and evaluate the effectiveness of each component in our adaptive watermark.
Entropy Thresholds. In Table 3, we provide the ROC-AUC, Best F1 Score under different entropy thresholds when the watermarked text is paraphrased using GPT-3.5. In addition, we provide the Average Watermark Rate (AWR), which is the fraction of the watermarked tokens. Figure 4 (left) shows the text perplexity comparison under different entropy thresholds. When $\alpha=0$, AWTI is skipped in watermark generation. In this case, the ROC-AUC and Best F1 Score are high, but the watermarked text exhibits high perplexity, with a median value exceeding $10$. When setting $\alpha$ to 1, 2, or 3, there is a notable decrease in the perplexity of the watermarked text, as fewer tokens are perturbed, reflected by the decreasing AWR. In addition, the robustness of the watermark is improved when AWTI is applied ($\alpha=1,2$), potentially because some modified tokens are filtered out by AWTI in detection. For entropy threshold $\alpha=3$, our method can achieve results comparable to other methods, with only about half of the watermarked tokens.
@@
Figure 4: Comparison of text perplexity at varying entropy thresholds (left) and across different watermark strengths (right). $\alpha$ represents the entropy threshold. $\delta$ represents the watermark strength.
Watermark Strength. In Table 4 and Figure 4 (right), we increase the watermark strength $\delta$ from $0.5$ to $2$ to examine its impact on robustness and perplexity. As shown in Table 4, there is little impact on the ROC-AUC values in terms of robustness. Interestingly, in contrast to previous works (Kirchenbauer et al., 2023a; Zhao et al., 2023) where perplexity tends to rise with larger watermark strength, the perplexity of the text watermarked with our method decreases as the $\delta$ increases in a reasonable range. This benefits from the proposed AWTS, which embeds the watermark by adjusting temperature as explained in Section 3.3. However, this does not imply that a larger $\delta$ consistently results in better text quality. A higher watermark strength can cause the token selection process to resemble greedy sampling more closely, potentially leading to the generation of repetitive content.
@@ 6 Conclusion
In this paper, we propose a holistic LLM watermark method to achieve the following properties: strong robustness, high security, high text quality, and accurate detection without the need to access the original language model and prompt. Our experiments show that, compared to baseline methods, our approach demonstrates better performance under different paraphrase attacks. Notably, our method has a negligible impact on text perplexity when compared to un-watermarked text. Additionally, we highlight the security of our approach by employing the spoofing attack. Specifically, the capacity of language models differs among various models. Therefore, the model owner may need to tune the entropy threshold and watermark strength carefully based on the performance of the language model to get better watermarking performance. Moreover, it is recommended for watermarked model owners to train their own small-sized measurement model based on the watermarked model for performance and security considerations. In future research, we aim to conduct a theoretical analysis to deepen the understanding of the performance characteristics of the watermarking approach and offer an optimal solution.
Impact Statement
Adding a watermark to the LLMs can significantly enhance transparency and accountability by making it easier to identify AI-generated content. Its impact extends broadly in the era of AI, including safeguarding intellectual property (IP), reducing the dissemination of misinformation, and mitigating the misuse of AI-generated content across educational and other domains. Furthermore, looking ahead to societal consequences, effective watermarking has the potential to bolster public trust in AI technologies by offering a clear means of distinguishing between human and AI-generated content.
Appendix A Semantic Mapping Model Implementation Details
@@
The Semantic Mapping Model is a neural network that starts and ends with linear layers and contains two residual blocks in the middle. For the loss function provided in Section 3.2, the weight of each term is set as 1. We train the model with the standard stochastic gradient descent with batch size 128.
@@
As discussed in Section 3.2 , we rescale Euclidean distance between two sentence embeddings $d=D(u(S),u(S^{\prime}))$ to a wider range. The original range $d\in[L,U]$, where $L=0$ and $U$ is the maximum distance. To enlarge this range, we apply a linear transformation $(T:\mathbb{R}\rightarrow\mathbb{R})$ to the original distance $d$, resulting in a new distance $d^{\prime}=T(d)\in[L^{\prime},U^{\prime}]$ where $L^{\prime}<L$ and $U^{\prime}>U$. The linear transformation is expressed as
$T(d)=\left(\frac{d-L}{U-L}\right)\cdot(U^{\prime}-L^{\prime})+L^{\prime}.$ (6)
In our experiment, we set the $L^{\prime}=-2$ and $U^{\prime}=4$.
Appendix B Additional Evaluation Results
B.1 Experiment Results on Additional LLM
To further showcase the performance of our watermarking algorithm, we conduct experiments on a more powerful LLM, Mistral-7B, using the C4 dataset. Moreover, we use a more powerful LLM, GPT-3, to calculate the perplexity. As presented in Table 5 and Figure 5, our watermarking method not only demonstrates superior robustness performance but also maintains lower text perplexity compared to other methods.
Table 5: The comparison of robustness performance across various watermarking methods, conducted on Mistral-7B with C4 dataset
Language Model Evaluation KGW-0 KGW-1 KGW-2 EXP-edit Ours
Mistral-7B ROC-AUC $0.952$ $0.854$ $0.792$ $0.872$ $0.961$
Best F1 Score $0.887$ $0.785$ $0.747$ $0.833$ $0.902$
Figure 5: The comparison of text perplexity across various watermarking methods, conducted on Mistral-7B with C4 dataset, with the perplexity calculated using GPT-3.
B.2 Experiment Results on Additional Dataset
Besides utilizing C4, which generates subsequent sentences from the first two as prompts, we also consider the long-form question-answering dataset ELI5 to reflect real-world LLM use cases. We conduct experiments with the ELI5 dataset using the OPT-6.7B model to assess robustness performance and text perplexity. In these experiments, watermarked texts were paraphrased by GPT-3, and text perplexity was also calculated using GPT-3. Table 6 indicates that our robustness performance is on par with KGW-0, while Figure 6 demonstrates that the text perplexity associated with our watermarking technique is significantly lower compared to other methods.
Table 6: The comparison of robustness performance across various watermarking methods, conducted on OPT-6.7 with ELI5 dataset.
Language Model Evaluation KGW-0 KGW-1 KGW-2 EXP-edit Ours
OPT-6.7 ROC-AUC $0.971$ $0.781$ $0.642$ $0.886$ $0.967$
Best F1 Score $0.913$ $0.733$ $0.668$ $0.857$ $0.899$
Figure 6: The comparison of text perplexity across various watermarking methods, conducted on OPT-6.7 with ELI5 dataset, with the perplexity calculated using GPT-3.
B.3 ROC Curves
@@
Figure 7: Comparisons of ROC curves of different watermark methods applied to various language models against two different paraphrasing attacks.
B.4 TPR at Fixed Low FPR
In practical applications, it is essential to monitor the TPR at a consistently low FPR to guarantee that un-watermarked texts are not mistakenly identified as watermarked. We showcase the TPR scores at fixed FPR thresholds of 1% and 10%, respectively, to demonstrate robustness performance under OPT-6.7B, as depicted in the Table 7.
Table 7: Robustness performance at fixed low FPR.
Language Model Methods TPR@1%FPR TPR@1%FPR ROC-AUC Best F1 Score
OPT-6.7B KGW-0 $0.670$ $0.896$ $0.964$ $0.904$
KGW-1 $0.198$ $0.653$ $0.874$ $0.815$
KGW-2 $0.179$ $0.455$ $0.769$ $0.728$
EXP-edit $0.257$ $0.731$ $0.881$ $0.801$
Ours $0.779$ $0.949$ $0.981$ $0.932$
B.5 Repetition Rate
To examine the repetition problem of watermarked text, we compute the different N-gram repetition rates for the watermarked text generated by different watermarking methods, setting N to 1, 2, and 3 separately. Table 8 shows that the repetition rate of our method is similar to that of the EXP-edit, which is a distortion-free watermarking technique.
Table 8: Repetition rate for watermarked text generated by different watermarking methods.
Language Model Methods 1-gram 2-gram 3-gram 4-gram
OPT-6.7B KGW-0 $0.219$ $0.046$ $0.019$ $0.012$
KGW-1 $0.199$ $0.037$ $0.013$ $0.007$
KGW-2 $0.199$ $0.036$ $0.015$ $0.010$
EXP-edit $0.239$ $0.062$ $0.026$ $0.018$
Ours $0.241$ $0.066$ $0.026$ $0.014$
Appendix C Comparison with Additional Baselines
We note that Lee et al. (2023) also discuss the strategy of perturbing only the high-entropy distributions during generation. However, their method requires access to both the original LLMs and prompts. Accessing the original prompt during the detection phase can be impractical, and employing the source watermarked LLM for detection is not only time-consuming but also incurs substantial computational costs.
Moreover, Liu et al. (2023b) propose a semantic-based watermarking method, which also relies on the original prompts for detection. While this method enhances the security of the watermark, there remains a security risk, as attackers might fix the semantics of prompts to access the corresponding green token list. Our algorithm has a more stable and robust Semantic Mapping Model to handle the perturbation of semantics after modification (e.g., paraphrase), further improving our watermark’s robustness. Moreover, our proposed AWTI further strengthens the security and improves the quality of watermarked text.
We conduct experiments to evaluate the robustness performance (after paraphrasing by GPT-3) and the perplexity of watermarked text using our algorithm (calculated by GPT-3), comparing it against both SWEET (Lee et al., 2023) and SIR (Liu et al., 2023b). All experiments are based on the OPT-6.7B model. The results of robustness performance are presented in Table 9, and the results of text perplexity are presented in Figure 8. It is worth noting that the comparison is not completely fair, as both SWEET and SIR are not model-agnostic. This implies that their detection processes involve using either the original prompt or the source watermarked LLM. The results show that we achieve better robustness performance and lower perplexity.
Table 9: Robustness performance comparison among Adaptive Watermark, SWEET, and SIR methods.
Language Model Methods ROC-AUC Best F1 Score
OPT-6.7 SWEET $0.847$ $0.787$
SIR $0.963$ $0.903$
Ours $0.981$ $0.932$
Figure 8: The comparison of text perplexity among Adaptive Watermark, SWEET, and SIR methods, conducted on OPT-6.7, with the perplexity being calculated using GPT-3.
Appendix D Security Discussion
The security of a watermarking algorithm here refers to the difficulty an attacker has in statistically analyzing and decrypting the watermarking rules using the watermarked texts with the aim of forging the watermarks. For previous approaches such as KGW-$k$, the rule of green/red list can be simply recovered by analyzing token frequency after a specific fixed token, as the green/red list is purely determined using $k$ previous tokens. For our semantic mapping model, the input sentence embedding space is continuous, and it depends on all the previously generated tokens, which requires the attackers to analyze token frequency based on fixed semantics. However, it is hard in practice, as even with the same prompt, the semantics of the generated text could vary widely. Consequently, for attackers aiming to identify the green token list associated with a particular semantic meaning, more effort or considerably larger sample size is required for our algorithm to conduct a meaningful analysis.
In general, we believe that a watermarking algorithm with more complex procedures will necessitate greater effort or a larger dataset for attackers to forge watermarking. To see this, we consider the examples of KGW-0 and KGW-$k$. KGW-0 employs a globally fixed green/red token list for generating watermarked text, making it the most straightforward to decrypt, i.e., simply analyzing overall token frequency. In contrast, KGW-$k$ enhances security by utilizing previously generated tokens to formulate the green/red token list, requiring attackers to analyze token frequency in relation to specific tokens, thereby complicating the decryption process. Our watermarking algorithm further advances the process by deriving the green/red token list from the semantics of previously generated text and adaptively applying it to tokens with high entropy. As the attackers cannot accurately identify watermarked tokens (high entropy tokens measured using the specific measurement model), this adds another layer of difficulty in statistically analyzing the token frequency, even within a specific semantic context. Therefore, the complexity of our watermarking algorithm makes it more challenging for attackers to decrypt the watermarking rules, thereby enhancing the security of the watermarking algorithm.
Specifically, the implementation of AWTI results in fewer green tokens (positive entries in logits scaling vector $v^{(t)}$) within the generated text compared to other methods like KGW-0/KGW-1/KGW-2. However, the power of our watermark will not be affected. As outlined in Section 4 of our detection process, we prioritize filtering out potential un-watermarked tokens ($U$) and focus on the proportion of green tokens ($\hat{G}$) among potential watermarked tokens ($W$), rather than the proportion of green tokens ($G$) among all tokens ($U\cup W$) in the generated text, as done in KGW-0/KGW-1/KGW-2. Consequently, while the overall ratio of green tokens to all tokens in the generated text ($\frac{|G|}{|U\cup W|}$) may drop, the proportion of green tokens among potential watermarked tokens ($\frac{|\hat{G}|}{|W|}$) might not necessarily decrease. To validate this, we analyzed paraphrased watermarked texts, calculating the average percentage of green tokens across 500 texts generated by KGW-0 and comparing it with the average percentage of green tokens ($\frac{|\hat{G}|}{|W|}$) across 500 texts produced by our watermarking algorithm. The findings, detailed in Table 10, reveal that although the overall percentage of green tokens in the generated text ($\frac{|G|}{|U\cup W|}$) stands at 0.499, the percentage of green tokens relative to potential watermarked tokens ($\frac{|\hat{G}|}{|W|}$) reaches 0.705, surpassing that of KGW-0. Furthermore, since the overall percentage of green tokens in the generated text is only about 0.5, this indicates that our generated texts do not exhibit a noticeable bias towards a specific group of tokens. This characteristic makes it more challenging for attackers to decrypt the watermarked texts by analyzing token frequency.
Table 10: Green token percentage for different methods.
Methods Green token percentage ROC-AUC
KGW-0 $0.649$ $0.964$
Ours ($\frac{|G|}{|U\cup W|}$) $0.499$-
Ours ($\frac{|\hat{G}|}{|W|}$) $0.706$ $0.981$
Appendix E Additional Ablation Study
@@
Measurement Models. We investigate how different Measurement Models (MMs) impact the robustness of our watermarks and the quality of the generated text. For this purpose, we employ MMs with different sizes and types, such as GPT2 (124M), GPT2-large (774M), OPT-125M, and OPT-350M, in our evaluation. For all models, we set the $\alpha$ and $\delta$ as $2$ and $1.5$, respectively. As shown in Table 11, there is only a minor influence on the robustness of the watermark when different MMs are adopted.
@@
However, the effect on text quality is more nuanced. As indicated in Figure 9, models like GPT2-large and OPT-350M will result in lower perplexity compared to the GPT2 and OPT-125M. One primary reason is that different MMs produce different entropy estimates, which results in different AWRs even with the same entropy threshold. Generally, a higher AWR implies a larger number of watermarked tokens, which can slightly impact text quality.
@@
Appendix F Watermarked Text Generation time
In this section, we compare the watermarked text generation time and the watermark detection time for different approaches. For each method, we generate and detect 100 instances of watermarked text, each containing roughly 200 tokens. Subsequently, we compute the average time taken for both the generation and detection of each instance of watermarked text. For our method, we set the value of the entropy threshold $\alpha=2$ and the watermark strength $\delta=1.5$. We evaluate the generation and detection time of the proposed adaptive watermarking with two different Measurement models.
@@ Appendix F Watermarked Text Generation time
Appendix G Watermarked Text Example
We present some examples of watermarked text using our proposed method. Table LABEL:watermarked_text_example shows the un-watermarked text and watermarked text under different entropy thresholds ($\alpha$). The watermarked texts are generated using the OPT-6.7B model with the watermark strength $\delta=1.5$. The watermarked tokens are highlighted in yellow. As the value of $\alpha$ increases, it is evident that the proportion of watermarked tokens noticeably decreases.
@@ Appendix G Watermarked Text Example
Prompt It’s not always easy to keep mentally focused during running. Focus requires discipline and a certain level of mental toughness, meaning that you have developed the capacity to perform to the best of your ability despite external conditions or internal distractions.
@@
Un-watermarked text While it’s not impossible to train for a specific event and keep mentally focused at all times, it’s challenging. Here are three things you can do to help yourself. Plan Ahead The best time to think about race strategy is during warm-ups before any race. Once your legs get moving, they will tell you what they need—you just have to listen to them! It’s also a good idea to write down in detail the training plan for your event and make sure that you understand it completely. Plan your warm-up so that there is an emphasis on building speed without injury risk. Get Out of Your Head There are lots of other things to focus your attention on besides running. If you’re running around town, take a walk break every quarter mile or so. This allows you to stretch, cool off and rest, while keeping your muscles engaged.
@@
Watermarked text ($\alpha=1$) The ability and desire for discipline are developed through practice. The first step in developing discipline during running, especially if a runner wants a higher level performance, begins with setting goals and working towards meeting these goals. Goal setting allows runners an avenue to create the mental toughness needed to achieve their best results during training runs or races. Setting realistic expectations will also allow runners to work within their abilities instead of trying to perform beyond what they are currently capable of. Setting realistic expectations can also serve as motivation for runners to continue training despite the obstacles that life has to offer during the course of training or race day. It can also serve as motivation to continue training even when injuries or other problems make training challenging or downright unpleasant. For runners who want to set realistic expectations for themselves during training or race day, the below checklist can serve as guidelines: Mentally prepared - Have enough sleep the night before the race or training runs Stress free environment - Remove negative people from the environment Eat right - Eat healthy meals the day before the race or training runs
@@
Watermarked text ($\alpha=2$) The best runners are often described in terms like, focused, determined, and driven. You may have read that somewhere on a running magazine cover, but did anyone actually write that down? I was recently asked if I knew any specific exercises to help people stay mentally focused during training runs. My response wasNo,” but the question got me thinking more about how to develop mental toughness for runners. Staying mentally prepared is an important part of the training process that can be just as important as physical preparation, but many runners don’t give it much thought. It’s easy to see how physical preparation is important for race day performance, but what exactly does it mean when you talk about mental preparation? Here are five things you need to do to stay mentally prepared while running: 1. Have realistic goals and expectations. Some people set unrealistic expectations for themselves or others regarding the number of miles they’ll run each week, their times for certain races, or even their overall goal for their training program.
@@
Watermarked text ($\alpha=3$) The best runners are focused on a specific task, which allows for a level playing field. The same applies in any aspect in which concentration and discipline are necessary. In order for a runner, cyclist, tennis star, golffer, etc. to perform their best, they must have the ability to stay focused at all times. To achieve this state of focus or mental toughness requires training (mental conditioning) and practice. You do not have to run every day for hours, but just like anything else it’s about putting in the time and practicing until you get better over time. So how does this apply to your life? The best runners know that running is an important part of their lives, as well as their training routine and their overall health. The best runners also understand what is required to keep them mentally strong during the race. They know how to train themselves so that they will always be ready and prepared for the next race, regardless of external conditions or the obstacles within themselves.